jack is thinking of two numbers. The sum of twice the first and three times the second is 9. the difference of the first number and the second number is 22. find the numbers.
Answer by ewatrrr(24785)

2x + 3y = 9
x - y = 22
|
2x+ 3y = 9
3x -3y = 66
5x = 75
x = 15 and y = -7 15+-+y+=+22
Checking
2x + 3y = 9
30- 21 = 9 CHECKS!